OpenStudy (cheesecakekitten):

WILL MEDAL AND FAN!!!!! Add [4 5] + [1 3] A. [9 4] B. [5 8] C. [4 15] D. [19]

OpenStudy (f_jayyy):

is the number [45] or [4 5]

OpenStudy (steve816):

i think it's matrix

OpenStudy (cheesecakekitten):

[4 5]

OpenStudy (steve816):

the numbers 4 and 5 probably

OpenStudy (f_jayyy):

i was just thinking it meant to get the absolute value and add them together but that would be 58 and thats not a choice

OpenStudy (cheesecakekitten):

I think the spaces mean multiply. 4 * 5 = 20 and 1 * 3 = 3 so 20 * 3 = 60 which is 4 * 15. So the answer must be C. [4 15]. But I don't know for sure if I'm supposed to multiply.

OpenStudy (cheesecakekitten):

@NetflixAndNo Hi Idk y I mentioned you it's just nice having you around

NetflixAndNo (netflixandno):

Aw hii

OpenStudy (cheesecakekitten):

hOI!

OpenStudy (cheesecakekitten):

I'M TEMMIE